### v1

# Motion blur; Blind deblurring; Blind motion deblurring; Deep learning

Abstract
Motion deblurring is one of the fundamental problems of computer vision and has received continuous attention. The variability in blur, both within and across images, imposes limitations on non-blind deblurring techniques that rely on estimating the blur kernel. As a response, blind motion deblurring has emerged, aiming to restore clear and detailed images without prior knowledge of the blur type, fueled by the advancements in deep learning methodologies. Despite strides in this field, a comprehensive synthesis of recent progress in deep learning-based blind motion deblurring is notably absent. This paper fills that gap by providing an exhaustive overview of the role of deep learning in blind motion deblurring, encompassing datasets, evaluation metrics, and methods developed over the last six years. Specifically, we first introduce the types of motion blur and the fundamental principles of deblurring. Next, we outline the shortcomings of traditional non-blind deblurring algorithms, emphasizing the advantages of employing deep learning techniques for deblurring tasks. Following this, we categorize and summarize existing blind motion deblurring methods based on different backbone networks, including convolutional neural networks, generative adversarial networks, recurrent neural networks, and Transformer networks. Subsequently, we elaborate not only on the fundamental principles of these different categories but also provide a comprehensive summary and comparison of their advantages and limitations. Qualitative and quantitative experimental results conducted on four widely used datasets further compare the performance of state-of-the-art methods. Finally, an analysis of present challenges and future pathways aims to guide and inspire researchers in the dynamic domain of image deblurring using deep learning techniques. This review aspires to serve as an invaluable reference for those immersed in this field of study. All collected models, benchmark datasets, source code links, and codes for evaluation have been made publicly available at https://github.com/VisionVerse/Blind-Motion-Deblurring-Survey
keywords:
Motion blur; Blind deblurring; Blind motion deblurring; Deep learning
[…]
1 Introduction
Motion blur is a common artifact in images caused by the relative movement between the camera and the scene during the exposure Burdziakowski (2020); Peng et al. (2021). This effect can cause object contours in the image to be blurred or stretched, thereby reducing the clarity and detail of the image and affecting various computer vision tasks such as autonomous driving Rengarajan et al. (2016); Li et al. (2023), object segmentation Rajagopalan et al. (2023); Zhou et al. (2023) and scene analysis Zhang and Bergholm (1997); Zhou et al. (2022). Understanding the origins of motion blur is critical to developing effective elimination strategies.
The blurry process, as a fundamental degradation model, can be depicted as the convolution between a clear image and a blur kernel Chen et al. (2021); Bai et al. (2019); Zhang et al. (2022):
$\mathbf{B}=\mathbf{S}\otimes\mathbf{k}+\mathbf{n},$ (1)
where $\mathbf{B}$ and $\mathbf{S}$ are the observed blurred image and sharp image, respectively. $\mathbf{k}$ and $\mathbf{n}$ represent the blur kernel and noise. $\otimes$ is the convolution operation. When $\mathbf{k}$ is known, the inverse process of this task is referred to as non-blind deblurring Arjomand Bigdeli et al. (2017); Sun et al. (2014); Zhang et al. (2017). However, $\mathbf{k}$ is often unknown when dealing with blurred images, leading to tasks known as blind deblurring Pan et al. (2016); Chen et al. (2019); Hu et al. (2014); Yan et al. (2017); Liu et al. (2019).
Motion blur can be divided into two types: global motion blur and local motion blur. As shown in Figure 1 (b), global motion blur is caused by the shaking of the camera Zhai et al. (2023), resulting in consistent blurring across the entire image. As shown in Figure 1 (c), local motion blur occurs due to relative motion between the target object and the static background, resulting in blur Zhang et al. (2022).
Figure 1: Sharp-blurry images.
Figure 2: An overview of traditional and deep learning methods for blind motion deblurring.
Deblurring represents an ill-posed problem that cannot be directly solved. The traditional blind deblurring methods Levin (2006); Yang and Ji (2019); Fergus et al. (2006) start with the blur degradation model, employing mathematical modeling or convolutional networks to predict the blur kernel. Then they utilize the estimated blur kernel to restore the image through deconvolution. These approaches rely on prior knowledge of both the image and the blur kernel to address the ill-posedness in deblurring. As shown in the traditional method on Figure 2, the commonly used prior algorithms are sparsity prior Krishnan et al. (2011); Xu et al. (2013); Levin et al. (2011), patch prior Sun et al. (2013); Guo and Ma (2018); Tang et al. (2018), channel prior Pan et al. (2017); Ge et al. (2021); Cai et al. (2020), gradient prior Levin et al. (2009); Pan et al. (2014); Shao et al. (2020) and so on. The quality of image deblurring is ensured by the priori algorithms. These traditional methods heavily rely on blur kernel estimation, and the effectiveness of the deblurring process directly depends on the accuracy of the kernel estimation. However, in complex real-world scenarios, the blur kernel in the blurring process tends to be complex, variable, and difficult to parameterize, causing the traditional methods to often perform inadequately.
In recent years, deep learning technology has developed rapidly and has been widely applied in image processing Zhang et al. (2012); Zhou et al. (2024). Deep learning-based methods have powerful feature learning and mapping capabilities, allowing them to learn complex patterns of blur removal from extensive datasets Gu et al. (2020). Consequently, significant progress has been made in image deblurringZhao et al. (2022); Purohit et al. (2021); Chen et al. (2022); Yang and Yamac (2022); Nimisha et al. (2017) As depicted in Figure 3, existing blind motion deblurring methods can be categorized based on different backbone networks, namely: convolutional neural networks (CNN) Mao et al. (2016), generative adversarial networks (GAN) Goodfellow et al. (2014); Wang et al. (2023), recurrent neural networks (RNN) Zhang et al. (2018), and Transformer networks Zhao et al. (2023). The deep learning methods can adaptively learn the blur features according to the training data, thus achieving end-to-end image deblurring. They can directly generate sharp images from blurred ones, which improves the effect of image deblurring. Compared to the traditional methods, the deep learning methods Li et al. (2021); Yuan et al. (2020); Noroozi et al. (2017) demonstrate improved robustness in real-world scenarios and possess broader applicability.
In this paper, we focus on the application of deep learning in blind motion deblurring. Overall, the main contributions of this paper are as follows:
  • In this review, we comprehensively overview the research progress of deep learning in blind motion deblurring, en-compassing the causes of motion blur, blurred image datasets, evaluation metrics for image quality, and methods developed over the past six years.
  • Specifically, we introduce a classification framework for blind motion deblurring tasks based on backbone networks, categorizing existing methods into four classes: CNN-based, RNN-based, GAN-based, and Transformer-based approaches.
  • Furthermore, for each type of blind motion deblurring method, we provide an in-depth summary of their fundamental principles, analyzing the advantages and limitations of each approach. Qualitative and quantitative experimental results on four widely used datasets further showcase the performance differences among different blind motion deblurring methods.
  • Finally, we conduct a prospective analysis and summary of current research challenges and potential future directions in blind motion deblurring. This review aims to offer researchers and practitioners in the field a comprehensive under-standing and a valuable reference for further advancements.
Figure 3: Overview of deep learning methods for blind motion deblurring.
The subsequent sections of this paper are structured as follows: We first summarized the classification and research status of image blind motion deblurring methods in Sec. 2. Then, We briefly introduced the public datasets and evaluation metrics in Sec. 3. Next, Sec. 4 describes the basic principles of different types of methods in detail, and summarizes and compares the advancements and shortcomings of different methods. Qualitative and quantitative experimental results on four common datasets further compare the performance of different methods. Finally, in Sec. 5 and Sec. 6, we provide a prospective analysis of current research challenges and potential future directions in blind motion deblurring.
2 Deep Learning Methods for Blind Motion Deblurring
As shown in Figure 3, the foundational framework of deep learning-based image deblurring methods can be divided into four categories: CNN-based, RNN-based, GAN-based, and Transformer-based image deblurring methods. We discuss these methods in the following sections.
2.1 CNN-based Blind Motion Deblurring Methods
CNN-based image blind deblurring methods can be divided into two categories according to the process of image deblurring. One is the early two-stage image deblurring network shown in Figure 4 (a), and the other is the end-to-end image deblurring methods with the better performance shown in Figure 4 (b).
Figure 4: Two CNN-based blind motion deblurring frameworks.
2.1.1 Two-stage CNN-based Blind Motion Deblurring Methods
Early image blind deblurring methods are mainly for deblurring a single blur kernel image. The process of image deblurring was divided into two stages, as shown in Figure 4 (a). The first stage is estimating the blur kernel through a neural network. Then, the estimated blur kernel is used for deconvolution or inverse filtering operations on the blurred image to achieve the effect of image deblurring.
Classical methods such as: Schuler et al. Schuler et al. (2015) trained a blind image deconvolution network to deblur the image. This network primarily consists of three parts: a feature extraction module, a blur kernel estimation module, and an image estimation module. The estimated blur kernel from the blur kernel estimation module is used to iteratively update the deblurred image. Sun et al. Sun et al. (2015) proposed to estimate the heterogeneous motion blur through the CNN model and then deconvolute the blurred images, which can effectively estimate the spatially varying motion kernel and achieve a better effect of removing motion blur. To address the issue of the non-uniform blur, Cronje et al. Cronje (2015) proposed a CNN-based image deblurring method. This method trained two CNN networks, one for estimating the vector of non-uniform motion and the other for image deconvolution. It accurately estimated non-uniform blur and enhanced the effectiveness of image deblurring. Ayan Chakrabarti et al. Chakrabarti (2016) trained a neural network for estimating the blur kernel of blur images. Using the estimated single global blur kernel, they used non-blind deconvolution to obtain deblurred images. Gong et al. Gong et al. (2017) trained a fully convolutional neural network for estimating image motion flow. Then they preformed non-blind deconvolution to recover clear images from the estimated motion flow. This method avoids the iterative process of underlying image priors. Xu et al. Xu et al. (2017) proposed a deep convolutional neural network to extract sharp edges from blur images for kernel estimation, which does not require any coarse-to-fine strategy or edge selection, thus greatly simplifying the kernel estimation and reducing the computational amount. Kaufman et al. Kaufman and Fattal (2020) analyze the blurring kernel by applying convolutional layers and pooling operations to extract features at multiple scales. Subsequently, the estimated kernel, along with the input image, is fed into a synthetic network to perform non-blind deblurring.
These two-stage image deblurring methods rely too much on the blur kernel estimation of the first stage, and the estimated blur kernel quality is directly related to the effect of image deblurring. In reality, the blur is mostly uneven, and the size and direction of the blurred image are also uncertain. So in the real scenes, this method can not remove complex real blur very well.
Figure 5: Different end-to-end CNN-based blind motion deblurring network architectures.
2.1.2 End-to-end CNN-based Blind Motion Deblurring Methods
The end-to-end image deblurring method generates a clear image directly from the input blurred image. It utilizes neural networks to learn complex feature mapping relationships for better restoration of clear images. End-to-end image deblurring algorithms have made significant progress since Xu et al. Xu et al. (2014) first implemented end-to-end motion blur image restoration using CNNs. As shown in Figure 5, CNN-based end-to-end image deblurring methods can be categorized into three types: multi-scale, multi-patch, and multi-temporal structures.
Multi-scale Structure: As shown in Figure 5 (a), the multi-scale model extracts multi-scale feature information from blurry images. Based on this extracted information, it progressively restores the image from coarse to fine levels. In order to better handle non-uniform blur in dynamic scenes, Nah et al. Nah et al. (2017)proposed a multi-scale convolutional neural network for end-to-end restoration of blurred images. Following a similar approach, many subsequent researchers have improved on the multi-scale strategy to obtain better results Purohit and Rajagopalan (2020); Tao et al. (2018); Vitoria et al. (2022); Hemanth and Latha (2020). However, its limitation lies in scaling the blurred image to a lower resolution, often resulting in the loss of edge information. Overlaying multi-scale input images onto sub-networks gradually enhances image clarity from lower-level to higher-level sub-networks, inevitably incurring higher computational costs. To reduce the spatio-temporal complexity of employing a multi-scale strategy network, Cho et al. Cho et al. (2021) devised a network architecture that starts from coarse to fine, proposing a multi-input multi-output U-Net network. This approach allows for a better execution of deblurring from coarse to fine.
Multi-patch Structure: As shown in Figure 5 (b), the multi-patch structure segments the blurred input into multiple patches for prediction. Inspired by spatial pyramid matching, Zhang et al. Zhang et al. (2019) proposed a deep hierarchical end-to-end CNN model called DMPHN. This model used multiple patches for a fine-to-coarse operation to handle blurry images. To remove spatially varying blur, Suin et al. Suin et al. (2020) proposed a patch hierarchical attention architecture. It significantly enhances deblurring effects through effective pixel-wise adaptation and feature attention. Building upon DMPHN, Zhang et al. Zhang et al. (2023) proposed a self-supervised event-guided deep hierarchical multi-patch network, which handles blurry images through a layered representation from fine to coarse. By stacking patch networks to increase network depth, this model enhanced deblurring performance. However, when segmenting input or features, discontinuities in contextual information may arise, potentially generating artifacts and leading to suboptimal image deblurring in dynamic scene contexts Zamir et al. (2021).
Multi-temporal Structure: As shown in Figure 5 (c), the multi-temporal structure gradually eliminates non-uniform blur over multiple iterations. The multi-temporal framework Park et al. (2020) was introduced to handle non-uniform blur and exhibited significant performance improvements. However, rigid progressive training and inference processes might not adapt well to images with varying degrees of blur in different regions. Subsequent researchers have proposed numerous time-series-based deblurring methods, learning from different temporal scales of motion blur to achieve enhanced deblurring effects Zhang et al. (2023a, b).
Figure 6: A spatial variant RNN image deblurring method.
2.2 RNN-based Blind Motion Deblurring Methods
Recurrent Neural Networks (RNNs) are commonly used for handling sequential data such as text, language, and time series, demonstrating exceptional performance when dealing with data that exhibit temporal or sequential patterns Liping et al. (2022). Due to their memory-based neural network architecture, RNNs have found applications in image processing as well. For instance, Zhang et al. Zhang et al. (2018) proposed an end-to-end spatially variant RNN network for deblurring in dynamic scenes. As illustrated in Figure 6, the weights of RNNs are learned by deep CNNs. They analyzed the relationship between spatially variant RNNs and the deconvolution process, demonstrating the ability of spatially variant RNNs to simulate the deblurring process. Models trained using the proposed RNNs are notably smaller and faster. To better utilize the properties of spatially varying RNNs, Ren et al. Ren et al. (2021) extend a unidirectionally connected one-dimensional spatially varying RNN to a two-dimensional spatial RNN with three-way connectivity. 2D RNNs can capture a larger receptive field and learn denser propagation between neighboring pixels compared to 1D RNNs, further improving deblurring performance. Tao et al. Tao et al. (2018) proposed a multi-scale deep neural network for single image deblurring based on convolutional long short-term memory (Conv LSTM), effectively utilizing previous frames and image features. Dongwon Park et al. Park et al. (2020) developed a multi-temporal recurrent neural network (MT-RNN) based on recursive feature mapping for blind single-image deblurring, progressively deblurring a single image in an iterative process. In practical applications, CNNs and other models are often combined with RNNs to compensate for certain limitations in image processing, such as sequence length restrictions and computational complexity.
Figure 7: GAN-based deblurring algorithm flow.
2.3 GAN-based Blind Motion Deblurring Methods
With GANs Goodfellow et al. (2014) showing promising results in computer vision tasks, they have also been applied to image deblurring tasks. The basic process of image deblurring based on GANs is illustrated in Figure 7. The generator learns to restore clear images from blurry ones, while the discriminator assesses whether the generated clear images are real, providing feedback to adjust the training of both the generator and discriminator.
Ramakrishnan et al. Ramakrishnan et al. (2017) proposed a depth filter based on GAN. For batter restoration performance, it combines global jump connections and dense structure. Additionally, the model omits the blur kernel estimation process, significantly reducing the testing time required for practical applications. Lin et al. Lin et al. (2019) proposed an adversarial blurred region mining and refinement method. This method can drive CGAN Mirza and Osindero (2014) to learn real sharp image distributions globally and locally for blind image deblurring. Based on GANs, O.Kupyn et al. Kupyn et al. (2018) proposed an end-to-end image deblurring network termed deblurGAN.DeblurGAN was trained based on WGAN Arjovsky et al. (2017) with gradient penalty Gulrajani et al. (2017) and perceptual loss Johnson et al. (2016), achieving notable deblurring effects. Based on DeblurGAN, O. Kupyn et al. Kupyn et al. (2019) proposed DeblurGAN-v2. This model incorporates a feature pyramid network within the generator. The discriminator employs a relativistic discriminator and uses least squares loss to assess the generated images on both global and local scales. Moreover, recognizing the impact of backbone network selection on deblurring quality and efficiency, DeblurGAN-v2 utilizes an embedded architecture that allows the use of different backbone networks for feature propagation. Compared to DeblurGAN, DeblurGAN-v2 demonstrates significant improvements in both performance and efficiency. Peng et al. Peng et al. (2023) also proposed the MobileNetDeblur-GAN (MND-GAN) deblurring algorithm based on the DeblurGAN algorithm. This algorithm utilizes MobileNet as a feature extractor to build a feature pyramid network. Additionally, it employs a multi-scale discriminator to jointly train the generator, reducing the computational cost of the generator while ensuring deblurring quality.
To further enhance image deblurring performance, Zhang et al. Zheng et al. (2019) proposed a multi-scale deblurring network based on GANs. The model first restores the edge information of the blurred image and then utilizes clear edges to guide a multi-scale network for image deblurring. Starting from the blurry images themselves, Zhang et al. Zhang et al. (2020) proposed an image deblurring approach that combines two GAN models: a BGAN subnetwork for learning the image blurring process to generate motion-blurred images and a DBGAN for learning the restoration of motion-blurred images. By learning how to blur images, this method can better remove blurriness.
Supervised learning relies on a large amount of training data, but in reality, there are few real blurred-clear image pairs available for collection. Madam et al. Nimisha et al. (2018) proposed an unsupervised motion image deblurring method based on GANs. The model employs GANs to learn the mapping relationship from blurred images to clear images. To prevent mode collapse and maintain consistency between deblurred and clear image gradients, the model incorporates a deblurring CNN module and a gradient module to constrain the solution space and guide the deblurring process. GAN-based methods may encounter issues during training such as model collapse and gradient disappearanceMescheder et al. (2018), potentially resulting in artifacts and loss of details in the reconstructed images. Building upon the cycle-consistent generative adversarial network (CycleGAN) Zhu et al. (2017), Wen et al. Wen et al. (2021) proposed an unsupervised image deblurring method using a multi-adversarial optimization CycleGAN to address artifacts in high-resolution image generation. This method iteratively generates clear high-resolution images through multiple adversarial constraints. Considering the significant role of edge restoration in the structure of deblurring blurry images, this method introduces a structural-aware strategy. It utilizes edge information to guide the network in deblurring, enhancing the network’s ability to recover image details. Zhao et al. Zhao et al. (2023) introduced a contrast-guided multi-attention CycleGAN unsupervised algorithm for motion deblurring. It improved the deblurring performance of unsupervised algorithms. Lu et al. Lu et al. (2019) proposed an unsupervised method for single image deblurring without pairing training images termed UID-GAN. The method introduces a disentangled framework to split the content and blur features of blurred images, which improves the deblurring performance.
Due to the current drawbacks of unsupervised learning, such as large model sizes, lengthy inference times, and strict constraints on image resolution and domains, Zhao et al. Zhao et al. (2022) proposed an unsupervised blind image deblurring model termed frequency-domain contrastive loss constrained lightweight CycleGAN (FCL-GAN). It designed lightweight domain transformation units and parameter-free frequency domain contrast units to ensure the algorithm’s "lightweight" and "real-time" operation, overcoming the limitations of the generally large models and complex inference in unsupervised methods. Additionally, this model enhanced deblurring performance by introducing frequency-domain contrastive learning. Considering the practical application of image deblurring algorithms in real-world scenarios, Liu et al. Liu et al. (2022) proposed a lightweight generative adversarial network termed Ghost-DeblurGAN. While this algorithm may not reach the advanced levels of deblurring, its advantage lies in its small model size and short inference time.
2.4 Transformer-based Blind Motion Deblurring Methods
The Transformer, renowned for its powerful attention mechanism, utilizes self-attention (SA) to capture long-range dependencies among patches in image sequences and adaptability to given input content Vaswani et al. (2017); Liu et al. (2021); Zhou et al. (2022); Raghu et al. (2021). When applied in image deblurring, it can achieve significant improvements in deblurring effectiveness.
Wang et al. Wang et al. (2022) proposed an efficient and effective Transformer-based image restoration architecture called U-former by embedding local enhancement window Transformer blocks into a U-shaped structure Ronneberger et al. (2015). They inserted learnable multi-scale restoration modulators into their decoder to enhance the quality of image restoration. The basic architecture is depicted in Figure 8. Although Uformer has shown promising performance in image restoration tasks, the heavy computation involved in self-attention restricts the application of the transformer within the encoding-decoding framework to limited depths. This inevitably affects the effectiveness of image restoration. Building upon U-former, Feng et al. Ji et al. (2021) proposed a deep encoder-decoder image restoration network called U2-former, with Transformer as the core operation. This model employs two nested U-shaped structures to facilitate feature mapping at different scales, resulting in an enhanced image deblurring effect. To optimize computational efficiency, the model utilizes a feature filtering mechanism to discard low-quality features, thereby reducing unnecessary computations.
Figure 8: The Transformer-based deblurring network architecture.
To effectively eliminate non-uniform blur, Tsai et al. Tsai et al. (2022) proposed an efficient Transformer-based image deblurring algorithm termed Stripformer. This method reweights image features by computing intra-band and inter-band attention to address blur of different sizes and orientations. It achieves notable deblurring results without requiring extensive training data, enhancing the removal of non-uniform blur. Zamir et al. Zamir et al. (2022) proposed a Transformer-based image restoration network termed Restormer. This method employs a multi-head attention mechanism for feature fusion to handle high-resolution images in image restoration tasks. Additionally, this method utilizes multi-scale local-global representation learning to effectively utilize distant contextual information. It simultaneously reduces computations while enhancing the effectiveness of image restoration. Kong et al. Kong et al. (2023) proposed an asymmetric network based on an encoder-decoder architecture. They designed an effective frequency-domain self-attention solver in the decoder, reducing spatial complexity and computational demands through element-wise operations. Additionally, they employed a discriminative frequency-based feedforward network (DFFN) to discern which low-frequency and high-frequency feature information should be retained for potential clear image restoration. Recently, Zhong et al. Zhong et al. (2023) proposed a Blur Interpolation Transformer (BiT) to reveal potential temporal correlations in blurry encodings. Based on multi-scale residual Swin Transformer blocks, they introduced dual-ended temporal supervision and a time-symmetric integration strategy to generate effective time-varying motion-rendered features, effectively generating clear image sequences from blurry inputs. Zou et al. Zou and Ma (2023) explored the significant roles of edges and textures in image deblurring and proposed a method called Edgeformer. This method utilizes edge information to enhance the correlation between pixels. By employing an edge-enhanced feed-forward network, it distinguishes between edge and texture features, resulting in improved image deblurring. To fully leverage both local and non-local features in images, Wu et al. Wu et al. (2023) introduced the hierarchical pyramid attention transformer (HPAT) network, based on a transformer-based multi-level network structure. It employs hierarchical attention fusion modules (HAFM) to extract global contextual information, enhancing the feature representation capability. Using a fusion feed-forward neural network (F3N) for feature fusion has improved the deblurring effectiveness. Liang et al. Liang et al. (2023) used deep features extracted from pre-trained visual transformers (ViT) Dosovitskiy et al. (2020) to encourage the recovered image to become clear without sacrificing quantitative performance. They also proposed local MAE perceptual loss and global distribution perceptual loss to guide image deblurring.
The existing image deblurring methods mainly focus on global deblurring, which may affect the clarity of locally blurred images. Li et al. Li et al. (2023) proposed a local dynamic motion deblurring transformation network (LDM-ViT) based on adaptive window pruning transformer blocks (AdaWPT). In order to focus the network attention on localized blur regions, AdaWPT removes unnecessary windows. Only the blurred regions are allowed to participate in the deblurring process, reducing the amount of unnecessary computation. Additionally, using labeled local blur image datasets to train the predictor can be able to remove local motion blur effectively. Yan et al. Yan et al. (2023) introduced a Transformer-based dynamic scene image deblurring network called SharpFormer. They designed a Local-preserving Transformer block (LTransformer) that learns local features while retaining global representations. To effectively apply LTransformer blocks to mid-resolution features, they used dynamic convolution blocks to adaptively handle input non-uniform blur, enhancing the quality of deblurred images.
Table 1: Representative benchmark datasets for evaluating single image deblurring algorithms.
Real/Syn. Dataset Generative approach Sizes Train/Valid/Test
Synthetic Köhler at al. Köhler et al. (2012) Motion trajectory captured by 6D camera 4 sharp+48 blur Not divided
GoPro Nah et al. (2017) averaging over consecutive clear frames 3,214 pairs 2,103 / 0 / 1,111
HIDE Shen et al. (2019) averaging over consecutive clear frames 8,422 pairs 6,397 / 0 / 2,025
Blur-DVS Jiang et al. (2020) averaging over consecutive clear frames 2178 pairs 1782/ 0 / 396
Real RealBlur Rim et al. (2020)
blurred images: Low shutter speed cameras
clear images: High shutter speed cameras
4,738 pairs 3,758 / 0 /980
RsBlur Rim et al. (2022)
blurred images: The long exposure camera
clear images: the short exposure camera
13,358 pairs 8,878 / 1,120 / 3,360
ReLoBlur Li et al. (2023)
blurred images: The long exposure camera
clear images: short exposure time camera
2,405 pairs 2,010 / 0 /395
2.5 Advantages and limitations of the four architectures.
In image deblurring, the four architectures CNN, RNN, GAN, and Transformer have their own advantages and limitations:
CNN is widely used in image processing to capture local features and spatial information Li et al. (2018); Liu et al. (2021). Trained with large-scale datasets, CNN-based deblurring methods perform high efficiency and generalization capabilities in Table 2. The CNN structure is simple and suitable for image denoising and deblurring tasks Tian et al. (2020). However CNN-based methods may be limited by a fixed-size receptive field and may not be effective for image-deblurring tasks that deal with global information or long-range dependencies. In order to solve the problem of a limited receptive field, the more widely used method is to use dilated convolution.
As shown in Table 2, RNN applies to sequence data and can capture temporal or sequential dependencies, which may be useful in certain image sequence deblurring scenarios. There may be gradient vanishing or explosion problems when dealing with long-term dependencies Rusch and Mishra (2021). Moreover, for image deblurring tasks, RNN is not well suited to capture spatial information Ren et al. (2021). Therefore, RNNs are generally combined with other structures to accomplish image deblurring tasks.
GAN improves the realism of image generation using adversarial training, which helps to generate more realistic deblurred images in Table 2. However, the training process may be unstable, requiring a balance between generator and discriminator training Cui et al. (2021). There may also be pattern crashes or non-convergence of training patterns.
As listed in Table 2, Transformer is able to capture global information and solve the remote spatial dependence problem Yan et al. (2023), and has advantages in processing for some image tasks that require long-distance dependence. However, the image deblurring task involves the processing of a large number of pixels and has a high computational cost.
3 Datasets and Evaluation Metrics
3.1 Datasets
Most neural network-based image deblurring methods typically require paired images for training. Existing blurry image datasets can be roughly divided into two categories: synthetic datasets and real datasets. Synthetic datasets include the Köhler dataset, Blur-DVS dataset, GoPro dataset, and HIDE dataset, among others. Real image datasets comprise the RealBlur dataset, RsBlur dataset, ReLoBlur dataset, and so on. As shown in Figure 9, these are example images from some commonly used datasets. Table 1 summarizes and outlines the basic information of select datasets. We additionally provide bar graphs to compare the size of the datasets, as shown in Figure 10.
Köhler et al. Dataset: To simulate real camera shake, Köhler et al. Köhler et al. (2012) used a Stewart platform with six degrees of freedom to replay the camera’s motion trajectory. During the replay of the motion trajectory, a series of clear images and long-exposure blurry images were captured. The dataset comprises 4 clear images and 48 corresponding blurry images for 12 motion trajectories.
Figure 9: Some example images from (a) Köhler dataset, (b) HIDE dataset, (c) RealBlur-J dataset, (d) GoPro dataset, (e) ReLoBlur dataset.
Blur-DVS Dataset: Created by Jiang et al. Jiang et al. (2020) using the DAVIS240C camera. The creation process involved capturing an image sequence with a slow-motion camera and then synthesizing 2,178 pairs of blurred and clear images by averaging adjacent 7 frames. The dataset consists of 1,782 pairs for training and 396 pairs for testing. Additionally, the dataset provides 740 real blurry images without paired clear images.
GoPro Dataset: Developed by Nah et al. Nah et al. (2017), this dataset is one of the most commonly used deep learning methods. The dataset’s motion-blurred images are generated by integrating multiple instantaneous clear images within a time interval. Using a GoPro Hero4 Black camera capturing clear images at 240fps, they averaged over varying time windows to create blurred images, with the clear image at the window’s center serving as the ground truth. It includes 3,214 pairs of clear and blurred images, divided into 2,103 pairs for training and 1,111 pairs for testing.
HIDE Dataset: Introduced by Shen et al. Shen et al. (2019), this dataset focuses on motion blur in pedestrian and street scenes, including camera shake and object movement. The dataset was created by capturing videos at 240 fps using a GoPro Hero camera and then averaging frames from these high-frame-rate videos. They averaged 11 consecutive frames to generate blurred images and kept the center frame as the clear image. It comprises 8,422 clear-blurred image pairs, divided into 6,397 pairs for training and 2,025 pairs for testing.
RealBlur Dataset: Rim et al. Rim et al. (2020) created the RealBlur dataset to train and test deep deblurring methods on real blurry images. The dataset consists of two subsets sharing the same image content. One subset, RealBlur-J, is generated from the camera’s raw images, while the other, RealBlur-R, is generated from JPEG images processed by the camera’s ISP. It contains a total of 9,476 image pairs, with each subset offering 4,556 pairs of blurry and ground truth clear images for 232 low-light static scenes. The blurry images in the dataset are affected by camera shake and were captured in low-light scenarios such as night streets and indoor rooms, covering common motion blur scenes.
RsBlur Dataset: Developed by Jaesung Rim et al. Rim et al. (2022), the RsBlur dataset provides real blurry images of various outdoor scenes, each paired with a series of 9 clear images. It comprises 13,358 real blurry images from 697 scenes. For analysis, the dataset is divided into training, validation, and testing sets, containing 8,878, 1,120, and 3,360 blurry images, respectively, with 465, 58, and 174 scenes in each set.
ReLoBlur Dataset: Designed by Li et al. Li et al. (2023), ReLoBlur uses a simultaneous multi-band shutter device to capture local motion-blurred images along with their corresponding clear images, followed by post-processing for correction. Unlike synthetic blur datasets, ReLoBlur authentically showcases the mixing effect between locally moving objects and backgrounds, including overexposed, dimly lit real-life scenes, and complex blurred edges. It contains common real-life scenes such as pedestrians, vehicles, family members, pets, balls, plants, and furniture, among others. The dataset publicly provides 2,010 pairs of training set images and 395 pairs for testing. Additionally, the authors annotated the areas of local motion blur.
Figure 10: Histogram of the number of samples in the datasets.
Table 2: A review of deep learning based deblurring methods, including publication information, overview, advantages and limitations.
Methods Pub. Category Overview Advantages Limitations
DeepDeblur Nah et al. (2017)
CVPR 2017
CNN
Removing complex motion blur from coarse to fine using a multi-scale recurrent network
End-to-end removal of complex motion blur without the need for kernel estimation
The mode has a large number of parameters,making training difficult
DMPHN Zhang et al. (2019)
CVPR 2019
CNN
Estimated by segmenting the image into multiple blocks and then removing blur at different scales
Overcoming the limitations of depth deblurring models on superposition depths
Feature segmentation may lead to discontinuous contextual information
PSS-NSC Gao et al. (2019)
CVPR 2019
CNN
Multi-scale CNN deblurring networks are analyzed and a parameter-selective sharing scheme is proposed
Selective parameter sharing can improve the learning efficiency and generalization of the model
The introduction of new network structures and hopping connection mechanisms may have increased computational costs
DGN Li et al. (2020)
TIP 2020
CNN
An effective depth-guided network consisting of an image deblurring branch and a depth refinement branch is proposed
The use of depth maps allows for a better understanding of scene changes and facilitates the use of depth information
Deblurring quality depends on the accuracy and quality of the depth map
MSCAN Wan et al. (2020)
TSCVT 2020
CNN
A multi-scale channel attention network is proposed for more robust feature representation and higher quality image recovery
Further taking advantage of multi-scale models, codec modules and residual block structures
The model has many parameters and is difficult to train
SDWNet Zou et al. (2021)
ICCV 2021
CNN
A network structure combining extended convolution and wavelet reconstruction is proposed for image deblurring
Dilated convolution results in a larger perceptual field and reduces the loss of texture information during sampling
Introduction of wavelet transform increases the complexity of the network
MIMOU-Net Cho et al. (2021)
ICCV 2021
CNN
Introducing asymmetric feature fusion for gradual recovery of clear images by coder-decoder
Handling multi-scale blurry with low computational complexity
Improvement of deblurring accuracy increases computational effort and time consuming
MPRNet Zamir et al. (2021)
CVPR 2021
CNN
Multi-stage architecture, injecting supervision at each stage to incrementally improve degraded inputs
Horizontal connections exist between feature processing blocks to avoid information loss
High dependence on computing power
MSSNet Kim et al. (2022)
ECCV 2022
CNN
A deep learning-based multi-scale stage network from coarse to fine is proposed
Multi-scale stage configurations and inter-scale information propagation contribute to the network’s deblurring performance
Hyperparameter adjustment is more complex
HINet Chen et al. (2021)
CVPRW 2022
CNN
Propose semi-instantiation to one piece using multi-stage architecture for image recovery task
HINet architecture generalizes well and has short inference time
Requires a large number of image pairs for training
BANet Tsai et al. (2022)
TIP 2022
CNN
Localization of blur direction and size by blurry perception for adaptive blur removal
Short inference time can well support subsequent real-time applications
Large amounts of diversity data are needed for training to adapt to dynamic scene blurring
IRNeXt Cui et al. (2023)
ICML 2023
CNN
Integration of multi-stage mechanisms into a u-shaped network to remove different sizes of blur in a coarse-to-fine manner
The local attention module improves restoration quality by weighting high-frequency information
The performance improvement of deblurring depends on the number of MSM branches
SVRNN Zhang et al. (2018)
CVPR 2018
RNN
End-to-end trainable spatially-variant RNN for dynamic scene deblurring
TThe pixel-by-pixel weights of the RNN are learned by the deep CNN, which helps to remove blurring
Need for a larger feeling field, dealing with large areas and spatial variations
SRN Tao et al. (2018)
CVPR 2018
RNN
Blur removal is achieved using a decoder with skip connections and parameter sharing across three scale
Sharing network weights at different scales, thus significantly reducing training complexity
Ignoring the variations between feature scales can easily saturate the network’s performance
MT-RNN Park et al. (2020)
ECCV 2019
RNN
Gradual removal of non-uniform blurring over multiple iterations using a self-looping MT structure
Contextual information is continuous, which facilitates the handling of blurring in dynamic scenes
Inflexible progressive training may not remove non-uniform blurring very well
DeblurGAN Kupyn et al. (2018)
CVPR 2018
GAN
Conditional GAN-based network for generating realistic deblurred images
Uniform and non-uniform motion blur images with less blurring can be recovered
Neglecting the nonlinearity characteristics of dynamic scenes blur, the network sensory field is small and homogeneous
DeblurGAN-V2 Kupyn et al. (2019)
ICCV 2019
GAN
Use feature pyramid networks and extensive backbone networks for better accuracy
Improved image restoration efficiency and enhanced real-time performance
Single network structure, which tends to lead to saturation of the network model
DBGAN Zhang et al. (2020)
CVPR 2020
GAN
Better deblurring by combining learning to generate blurred images and deblurring images
Synthetic blurred images more accurately mimic real blur for better deblurring
Network training is complex
CycleGAN Zhu et al. (2017)
ICCV 2021
GAN
Blind motion deblurring based on multi-adversarial mechanism for iterative generation of high-resolution images
Add multi-scale edge constraint function to enhance network deblurring and detail preservation capability
Presence of large color differences, severe artifacts and model redundancy
FCL-GAN Zhao et al. (2022)
MM 2022
GAN
Constraining Lightweight Network CycleGAN Driven Tasks by Frequency Domain Contrast Losses
Better deblurring of images in a lightweight and real-time manner
Deblurring effect does not reach the supervised model SOTA level
Ghost-DeblurGAN Liu et al. (2022)
IROS 2022
GAN
Introducing GhostNet for designing lightweight deblurring networks
Suitable for real-time image deblurring, small model size, fast inference time
Deblurring effect fails to reach SOTA level
Uformer Wang et al. (2022)
CVPR 2021
Transformer
Building hierarchical codec networks with the Transformer module
Ability to capture local and global dependencies
Self-attentive layers can only be applied to a limited depth
Restormer Zamir et al. (2022)
CVPR 2022
Transformer
Codecs for learning multi-scale local-global representations of high-resolution images
Ability to model global connectivity with high computational efficiency
Image recovery does not fully utilize spatial information
Stripformer Tsai et al. (2022)
ECCV 2022
Transformer
Re-weighting image features horizontally and vertically to capture blurred features in different directions
Low memory footprint and computational cost, does not rely on large amounts of training data
Scale dot products Note that complex matrix multiplication is usually required
FFTformer Kong et al. (2023)
CVPR 2023
Transformer
Combining a self-attentive solver with a feed-forward neural network based on discriminative frequency domain to remove blur
Lower space and time complexity and more effective and efficient in image deblurring
May not generalize well to different types of images and blurs
BiT Zhong et al. (2023)
CVPR 2023
Transformer
Introducing Double-Ended Temporal Supervision and Temporal Symmetry Integration Strategies to Generate Effective Time-Varying Motion Rendering Features
Capable of effectively handling blurring at different scales and fusing information from neighboring frames
Limited discrete supervision does not respond well to different realities
Sharpformer Yan et al. (2023)
TIP 2023
Transformer
Adaptive handling of non-uniform blurring with dynamic convolutional blocks by fully utilizing global-local features
Able to consistently improve the quality of deblurring due to CNN-based deblurring methods alone
Higher computational complexity
3.2 Evaluation Metrics
Image quality assessment (IQA) is an important method to evaluate the effectiveness of the model. Current image quality evaluation methods can be divided into subjective and objective evaluation indexes.
3.2.1 Subjective Evaluation
Subjective evaluation relies on human judgment and is often the most intuitive method for assessing image quality. One representative measure is the Mean Opinion Score (MOS) Hoßfeld et al. (2016). Subjective evaluation scales primarily include absolute and relative scales. As shown in Table 3, individuals rate image quality within the range of 1-5. However, when different deblurring methods produce similar results, subjective evaluation often struggles to differentiate subtle differences. This susceptibility to individual biases and external environmental factors during assessment can lead to varying subjective evaluation results among different observers. Therefore, subjective evaluation is generally used as a supplementary assessment method in image quality evaluation and cannot entirely determine the final image quality assessment outcome. As a result, for image deblurring, most existing methods rely on objective evaluation metrics for assessment.
Table 3: Subjective evaluation scale.
Score normal-$\uparrow$ Quality scale Relative quality Absolute quality
5 There is no sign of the image quality deteriorating. The best. Beyond compare
4 The image quality is poor, but it is still viewable. Better than the average level. Nice
3 The image quality is poor, slightly hampers viewing. The average level. Commonly
2 Hindrance to watching. Below average in favorability. Bad
1 Very serious obstruction of viewing. The worst Very bad
Table 4: Performance comparison of state-of-the-art methods on four commonly used image deblurring datasets: GoPro, HIDE, RealBlur-J, and RealBlur-R. In red are depicted the overall best results among all methods, while in bold are showcased the best results within each respective category.
Methods Category GoPro Nah et al. (2017) HIDE Shen et al. (2019) RealBlur-J Rim et al. (2020) Realblur-R Rim et al. (2020)
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
DeepDeblur Nah et al. (2017) CNN 29.08 0.914 25.73 0.874 27.87 0.834 32.51 0.841
DMPHN Zhang et al. (2019) CNN 31.20 0.940 29.09 0.924 28.42 0.860 35.70 0.948
SDWNet Zou et al. (2021) CNN 31.26 0.966 28.99 0.957 28.61 0.867 35.85 0.948
MSCAN Wan et al. (2020) CNN 31.24 0.945 29.63 0.920 - - - -
PSS-NSC Gao et al. (2019) CNN 31.58 0.948 - - - - - -
MIMOU-Net+ Cho et al. (2021) CNN 32.45 0.957 29.99 0.930 27.63 0.837 35.54 0.947
MPRNet Zamir et al. (2021) CNN 32.66 0.959 30.96 0.939 28.70 0.873 35.99 0.952
HINet Chen et al. (2021) CNN 32.77 0.959 - - - - - -
BANet Tsai et al. (2022) CNN 32.54 0.957 30.16 0.930 - - - -
SAPHNet Suin et al. (2020) CNN 32.02 0.953 29.98 0.930 - - - -
MSSNet Kim et al. (2022) CNN 33.01 0.961 30.79 0.939 28.79 0.879 35.93 0.953
LaKDNet Ruan et al. (2023) CNN 33.35 0.964 31.21 0.943 28.78 0.878 35.91 0.954
SRN Tao et al. (2018) RNN 30.26 0.934 28.36 0.915 28.56 0.867 35.66 0.947
MT-RNN Park et al. (2020) RNN 31.15 0.945 29.15 0.918 28.44 0.862 35.79 0.951
ESTRNN Zhong et al. (2023) RNN 31.07 0.902 - - - - - -
IFI-RNN Nah et al. (2019) RNN 29.97 0.895 - - - - - -
RNN-MBP Zhu et al. (2022) RNN 33.32 0.963 - - - - - -
DeblurGAN Kupyn et al. (2018) GAN 28.70 0.858 24.51 0.871 27.97 0.834 33.79 0.903
DeblurGAN-V2 Kupyn et al. (2019) GAN 29.55 0.934 26.61 0.875 28.70 0.866 35.26 0.944
DBGAN Zhang et al. (2020) GAN 31.10 0.942 28.94 0.915 24.93 0.745 33.78 0.909
CycleGAN Zhu et al. (2017) GAN 22.54 0.720 21.81 0.690 19.79 0.633 12.38 0.242
FCL-GAN Zhao et al. (2022) GAN 24.84 0.771 23.43 0.732 25.35 0.736 28.37 0.663
Ghost-DeblurGAN Liu et al. (2022) GAN 28.75 0.919 - - - - - -
CTMS Zhao et al. (2022) Transformer 32.73 0.959 31.05 0.940 27.18 0.883 - -
Uformer Wang et al. (2022) Transformer 32.97 0.967 30.83 0.952 29.06 0.884 36.22 0.957
Restormer Zamir et al. (2022) Transformer 32.92 0.961 31.22 0.942 28.96 0.879 36.19 0.957
Stripformer Tsai et al. (2022) Transformer 33.08 0.962 31.03 0.940 - - - -
Stoformer Xiao et al. (2022) Transformer 33.24 0.964 30.99 0.941 - - - -
FFTformer Kong et al. (2023) Transformer 34.21 0.969 31.62 0.945 - - - -
Sharpformer Yan et al. (2023) Transformer 33.10 0.963 - - - - - -
3.2.2 Objective Evaluation
Objective assessment primarily involves quantitatively analyzing and evaluating deblurred images through the construction of mathematical models Su et al. (2022); Li (2022); Wang et al. (2004). Objective assessment enables the unified measurement of various deblurring methods through precise data calculations, offering advantages of fairness, objectivity, and efficiency. Among these evaluation methods, Mean Squared Error (MSE), Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM) stand as the most commonly employed metrics in image restoration tasks.
Mean Square Error (MSE). MSE is the average of the squared absolute errors between the restored image and the ground-truth sharp image. The closer the MSE is to zero, the smaller the difference between the restored image and the ground truth image, indicating a more realistic restoration Li (2022).
$\operatorname{MSE}(\mathbf{I},\mathbf{S})=\frac{1}{K}\sum_{k=0}^{K-1}\left(\mathbf{I}_{k}-\mathbf{S}_{k}\right)^{2},$ (2)
where $\mathbf{I}$ and $\mathbf{S}$ represent the reconstructed image and the real sharp image, respectively. $K$ denotes all pixels in the image.
Peak Signal Noise Ratio (PSNR). PSNR Hore and Ziou (2010) is a commonly used metric to assess the effectiveness of image deblurring. It measures the accuracy of image deblurring by comparing the MSE between the deblurred and sharp images. However, PSNR only focuses on differences at the pixel level. In image deblurring assessment, it’s typical to combine other metrics to comprehensively evaluate image quality. PSNR is defined as:
$\operatorname{PSNR}(\mathbf{I},\mathbf{S})=10\log_{10}\left(\frac{\left(2^{n}-1\right)^{2}}{\operatorname{MSE(\mathbf{I},\mathbf{S})}}\right),$ (3)
where $n$ is the number of digits where the image is stored. The larger the PSNR value is, the closer the recovered image is to the original image, the better the image recovery is.
Structural Similarity Index Measure(SSIM). SSIM Wang et al. (2004) is one of the commonly used metrics in image quality assessment. In the task of image deblurring, it evaluates the similarity between the deblurred image and the ground-truth sharp image by comparing their structural information. SSIM primarily assesses image similarity based on three characteristics: luminance, contrast, and structure. It doesn’t just focus on pixel differences but also considers features like image structure and texture. Compared to PSNR, SSIM aligns better with human visual perception. The calculation formula is as follows:
$\operatorname{SSIM}(\mathbf{I},\mathbf{S})=[l(\mathbf{I},\mathbf{S})]^{\alpha}\cdot[c(\mathbf{I},\mathbf{S})]^{\beta}\cdot[s(\mathbf{I},\mathbf{S})]^{\gamma},$ (4)
where $l(\mathbf{I},\mathbf{S})$ represents the brightness comparison of the image, $c(\mathbf{I},\mathbf{S})$ represents the contrast comparison of the image, $s(\mathbf{I},\mathbf{S})$ represents the structural comparison of the image. In engineering, often make $\alpha=\beta=\gamma=1$. In Eq. (4),
$l(\mathbf{I},\mathbf{S})=\frac{2\mu_{I}\mu_{S}+C_{1}}{\mu_{I}^{2}+\mu_{S}^{2}+C_{1}},~{}~{}~{}c(\mathbf{I},\mathbf{S})=\frac{2\sigma_{I}\sigma_{S}+C_{2}}{\sigma_{I}^{2}+\sigma_{S}^{2}+C_{2}},~{}~{}~{}s(\mathbf{I},\mathbf{S})=\frac{\sigma_{IS}+C_{3}}{\sigma_{I}\sigma_{S}+C_{3}},$ (5)
where $\mu_{I}$, $\sigma_{I}$ and $\mu_{S}$, $\sigma_{S}$ represent the mean, standard deviation of the images $\mathbf{I}$ and $\mathbf{S}$, respectively. $\sigma_{IS}=\frac{1}{K}\sum_{k=0}^{K-1}\left(\mathbf{I}_{k}-\mu_{I}\right)\left(\mathbf{S}_{k}-\mu_{S}\right)$ indicates the covariance of the reconstructed image $\mathbf{I}$ and the real sharp image $\mathbf{S}$. $C$ is a constant to avoid the denominator 0 causing system errors. Both $C_{1}$, $C_{2}$ and $C_{3}$ are constants, with values ranging from $[-1,1]$. A higher SSIM value indicates a greater similarity between the restored image and the original clear image, signifying a better restoration outcome.
Figure 11: Performance comparison (PSNR, %) of SOTA methods on the GoPro dataset.
Table 5: Speed and Parameters of representative SOTA methods.
Method Category Param. (M) Time (s) PSNR (%)
DMPHN Zhang et al. (2019)
CNN
21.7
0.21
31.20
MIMOU-Net+ Cho et al. (2021)
CNN
16.1
0.02
32.45
MPRNet Zamir et al. (2021)
CNN
23.0
0.09
32.66
SRN Tao et al. (2018)
RNN
6.8
0.07
30.26
DeblurGAN-V2 Kupyn et al. (2019)
GAN
60.9
0.04
29.55
Restormer Zamir et al. (2022)
Transformer
26.1
0.08
32.92
Uformer Wang et al. (2022)
Transformer
50.9
0.07
32.97
Stripformer Tsai et al. (2022)
Transformer
19.7
0.04
33.08
FFTformer Kong et al. (2023)
Transformer
16.6
0.13
34.21
Figure 12: Comparison of background deblurring results on the GoPro dataset.
Figure 13: Comparison of human deblurring results on the GoPro dataset.
4 Comparison with State-of-the-Art Methods
4.1 Quantitative Comparison.
Quantitative comparisons were conducted among 30 blind motion deblurring methods across four datasets, encompassing 12 CNN-based, 5 RNN-based, 6 GAN-based, and 7 Transformer-based approaches. From Table 4, it is evident that Transformer-based methods exhibit competitive performance across the majority of datasets. This can be attributed to the self-attention mechanism in Transformers, which adeptly captures remote dependencies between local and global image sequences, enhancing the model’s ability to recover details from blurred images. As depicted in Table 4, minimal performance disparities were observed between CNN and RNN methods on the GoPro dataset, both achieving over 33% PSNR and 0.96 SSIM. This similarity can be attributed partly to their analogous structures and the relatively simple motion blur in the synthetic nature of the GoPro dataset. Figure 11 illustrates that FFTformer achieved the highest PSNR score at 34.21%. Conversely, as indicated in Figure 11, the performance of GAN-based methods was inferior to other types due primarily to challenges related to training difficulties and convergence issues within GAN models, resulting in suboptimal deblurring restoration effects.
In addition, Table 5 presents a comparative analysis of various methods, balancing parameter count and operational speed. It is evident from Table 5 that RNN-based methods exhibit the smallest parameter count, CNN-based methods demonstrate the fastest operational speed, while Transformer-based methods yield the best performance. This insight further guides us toward integrating the strengths of different methodological types in subsequent research endeavors, aiming to achieve high-performance, real-time blind motion deblurring methods.
4.2 Qualitative Comparison.
As shown in Table 2, this review summarizes over 30 state-of-the-art blind motion deblurring methods based on deep learning. It includes publication details, summaries, advantages, and limitations. Among these, 22 were published in top-tier computer vision conferences such as CVPR, and ICCV, while 8 were featured in leading image processing journals like TIP, and TCSVT, among others.
To visually compare the deblurring performance of different algorithms, we conducted a qualitative assessment of selected algorithmic performance on the GoPro dataset. Figure 12 and Figure 13 illustrate the deblurring effects of various methods on backgrounds and human subjects, respectively. DBGAN and Uformer depict the deblurring results published by the original authors, while the remaining deblurred results were reproduced by us. The experiments were conducted on a Windows 10 system using Python 3.8 and Torch 2.0.0. As depicted in Figure 12 and Figure 13, the deep learning-based blind image deblurring methods tested exhibit effective image restoration. While some differences persist compared to the original clear images, visually, they do not significantly impact the image information. Combining the insights from Table 4 and Figure 13, it becomes apparent that Transformer-based models not only outperform other methods in evaluation metrics such as PSNR and SSIM but also demonstrate superior restoration of image details throughout the image recovery process.
Table 6: Comparison of generalization performance on real motion blur datasets.
Methods Trained on GoPro Trained on Realblur
RealBlur-J Realblur-R RealBlur-J Realblur-R
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
SRN Tao et al. (2018) 28.56 0.867 35.66 0.947 31.38 0.909 38.65 0.965
SDWNet Zou et al. (2021) 28.61 0.867 35.85 0.948 30.73 0.896 38.21 0.963
DeblurGAN-V2 Kupyn et al. (2019) 28.70 0.866 35.26 0.944 29.69 0.870 36.44 0.935
MPRNet Zamir et al. (2021) 28.70 0.873 35.99 0.952 31.76 0.922 39.31 0.972
MSSNet Kim et al. (2022) 28.79 0.879 35.93 0.953 32.10 0.928 39.76 0.972
LaKDNet Ruan et al. (2023) 28.78 0.878 35.91 0.954 32.33 0.929 39.91 0.974
MAXIM-3S Tu et al. (2022) 28.83 0.875 35.78 0.947 32.84 0.935 39.45 0.962
DeepRFT+ Mao et al. (2023) 28.97 0.884 36.06 0.954 32.19 0.931 39.84 0.972
5 Current Challenges and Future Prospects
5.1 Challenges of Blind Motion Deblurring and Major Concerns
The paper provides an overview of image deblurring algorithms based on deep learning, encompassing methods such as CNN-based, RNN-based, GAN-based, and Transformer-based image deblurring techniques. These methods are analyzed, compared, and summarized, and their performance is contrasted.
Despite the progress made in image deblurring algorithms concerning restoration effects, challenges persist due to the complexity of real-world blur scenarios. The challenges of blind motion deblurring and major concerns are as follows:
(1) Generalization Ability of Model: As shown in Table 6, the model trained on synthetic datasets exhibits poor generalization on real datasets. Compared with the training on the real datasets, the overall performance of the methods trained on the synthetic dataset decreases by about 3%, and the SSIM value will decrease by at least 0.02 in Table 6. Enhancing the generalization of model is a question that needs to be addressed.
(2) Data Requirements and Diversity: Most deep learning-based image deblurring algorithms rely on training with clear-blurred image pairs. The types of motion blur are diverse. Handling different types and intensities of motion blur requires a large amount of diverse data. Collecting and labeling this data might be a challenge, especially considering the types of blur, lighting conditions, and scene variations.
(3) Balancing Model Performance and Light-weight Models: Increasing model depth and parameters can enhance fitting data and improve image restoration quality. However, this can lead to larger network sizes, longer inference times, and higher computational complexity. Finding a balance between model performance and model light-weighting for practical applications is an essential area for exploration.
(4) Deblurring Methods for High-Level Visual Tasks: Blurred images hinder several advanced computer vision tasks, yet many existing deblurring methods aim solely at recovering high-quality images without considering aiding higher-level tasks such as object detection Zheng et al. (2021); Wu et al. (2023); Zhao et al. (2023) or image segmentation Pan et al. (2016); Luo et al. (2021). Research focusing on effectively extracting relevant features for high-level visual problems to enhance image deblurring solutions is a promising direction.
5.2 Future Prospects
(1) More comprehensive datasets: Most deep learning-based image deblurring algorithms rely on training data comprising blurred images. Constructing a more comprehensive dataset that encompasses a broader range of application scenarios can empower models to better handle various real-world situations, such as motion blur, low-light conditions, and lens shake. However, building such a comprehensive dataset poses challenges. Collecting and labeling various types of blurry images may require substantial time and resources. Moreover, ensuring the quality and diversity of the dataset is crucial, as low-quality or imbalanced data could adversely affect model training.
(2) More innovative network structure: Considering the complexity of image deblurring and the practical demands of applications, it’s essential to innovate and improve the deblurring network architectures. Combining the latest advancements in deep learning, these architectures may better capture complex blur patterns within images. This enhancement could enable them to address various types and intensities of blur, allowing models to perform exceptionally well across a broader range of scenarios.
(3) More objective evaluation indicators: Current evaluation metrics may fail to capture all subtle variations in the image deblurring process. Hence, more objective and comprehensive metrics are crucial for assessing algorithm effectiveness and performance enhancements. Considering multiple aspects such as clarity, detail retention, noise, contrast, etc., constructing comprehensive evaluation metrics would better assess the effectiveness of image deblurring. Introducing evaluation metrics aligned with human visual perception, beyond pixel-level metrics, considering the human perception of details, contrast, color, etc., might better reflect the performance of image deblurring techniques.
(4) Few-shot learning: Traditional deep learning models often require extensive labeled data for training, which might be challenging to obtain, especially high-quality labeled clear images. Few-shot learning Zhang et al. (2019); Chi et al. (2021); Yang et al. (2022) help models rapidly learn and adapt to new tasks from a limited amount of labeled data, reducing dependence on extensive labeled datasets. This approach addresses challenges related to acquiring and labeling data, potentially widening the application of image deblurring techniques across real-world scenarios.
(5) Applications of diffusion: The Diffusion model is a generative model that has garnered significant attention in recent years, demonstrating immense potential in the realm of image editing and enhancement. In the field of image deblurring, the Diffusion model has also found applications  Ren et al. (2023); Whang et al. (2022). Its flexibility allows it to handle various types of noise and transformations, suggesting its capability to potentially address complex image blurring. In the future, with a deeper understanding of the Diffusion model and the development of related techniques, it is poised to become a significant tool in the domain of image deblurring, promising more satisfying results and improvements.
(6) Unsupervised Deblurring Methods: Real-world blurry images often lack corresponding clear counterparts. Employing unsupervised methods Zhao et al. (2022) could render algorithms robust against real-world blurry scenarios, operating without extensive training data.
6 Conclusions
The current research and developments indicate significant progress in enhancing image quality and clarity through deep learning technologies. This paper reviews recent advancements in deep learning-based image deblurring algorithms, providing a comprehensive classification and analysis. It elaborates on existing image deblurring algorithms from four perspectives: CNN, RNN, GAN, and Transformer. Additionally, it conducts comparative analyses of representative algorithms, summarizing their strengths and limitations. Finally, the paper outlines future research directions in image deblurring algorithms. However, it is acknowledged that addressing more complex blurs in real-world scenarios remains a challenge. The quality of data and the accuracy of labels are crucial for training deep learning models, necessitating large-scale, high-quality datasets to support model training and optimization. In the future, there is an anticipation to further optimize deep learning models to enhance their speed and efficiency, expanding the application of these techniques in fields such as video processing, surveillance, and autonomous driving.
[…]
Methodology, Y.X.; Supervision, Z.L. and Y.X.; Writing—Original Draft, Y.X., H.Z. and C.L.; Writing—Review and Editing, Y.X., H.Z. and F.S.; Visualization, Y.X., F.S., C.L. and H.Z.; All authors have read and agreed to the published version of the manuscript.
[…]
This research received no external funding.
[…]
Not applicable.
[…]
The authors declare no conflict of interest.